OpenStudy (littlebird):

find a parametrization for the left half of the parabola y=x^2-4x+3 ?

OpenStudy (anonymous):

hen you substitute your x(t) function for into; y(x(t)) = x(t)^2 - 4x(t) + 3 but since x(t) = t, then y(t) = t^2 - 4t +3 Hope that makes sense, W.O.M.B.A.T.

OpenStudy (littlebird):

You copied that from yahoo. I don't understand why they did what they did in that solution.